OpenStudy (anonymous):

find k such that the line kx+y=3,is parallel to the line 5x-2y=1

OpenStudy (saifoo.khan):

Note, k is the coefficent of x... so k is the slopel

OpenStudy (anonymous):

ok go on

OpenStudy (anonymous):

hv tried but didn't gget the answer,pls help me solve it

OpenStudy (saifoo.khan):

5x-2y=1 5x - 1 = 2y 2y = 5x - 1 y = 5x/2 -1/2 So slope = 5/2

OpenStudy (saifoo.khan):

Since that is parallel, so the value for k is 5.2

OpenStudy (saifoo.khan):

5/2

OpenStudy (anonymous):

thanks dear